OpenStudy (anonymous):

Let c be any constant number. Which of the following will always be perpendicular to −3x + y = 2? A: y=4x+c B: y=-4+c C: y=-1/4x+c D: y=1/4x+c

OpenStudy (anonymous):

to find if two linear lines are perpendicular you would need to find a the slope of one line =m, then find the slope of another line that is 1/m.

OpenStudy (anonymous):

i don't understand how todo that @Chris9k

OpenStudy (anonymous):

Do you know y = mx + b?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

That is the equation of a line. all of your answers are in that form. If you take your first equation.. -3x + y =2 and make it match the others..y = 3x +2 now compare that to y = mx + b....what is m for that equation?

OpenStudy (anonymous):

-3x?

OpenStudy (anonymous):

the x is already there..just the m...and look at the second equation I made...y = 3x + 2

OpenStudy (anonymous):

im confused. :(

OpenStudy (anonymous):

Okay..let's start over. The equation of a line is y= mx + b m is the slope ok?

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

The equation they gave you is -3x +y = 2 How would you make that y = ? By adding 3x to both sides. right?

OpenStudy (anonymous):